OpenStudy (anonymous):

Rewrite y = x2 + 4x + 5 in vertex form. Then find the vertex.

OpenStudy (mchilds15):

It's in standard form right now.... do you know vertex form?

OpenStudy (zale101):

\(\Large x^2 + 4x + 5=y\) subtract 5 to both sides \(\Large x^2 + 4x =y-5\) take half of the coefficent and square, the coffeicent is 4x \((\frac{4}{2})^2\)

OpenStudy (mchilds15):

-.- zale, you take the fun out of it

OpenStudy (zale101):

My bad :|

OpenStudy (zale101):

@bg5150 are you working on it? :o

OpenStudy (anonymous):

sorry guys i had to run to the jon haha im gonna get on it

OpenStudy (zale101):

what does the vertex form look like?

OpenStudy (anonymous):

is it the one you just posted?

OpenStudy (zale101):

No, vertex form is \(\Large y=(x-\color{red}h)^2+\color{blue} k\)

OpenStudy (zale101):

where h,k is the vertex of the parabola

OpenStudy (zale101):

what i suggested above was just the pathway of how to get the equation converted in vertex form.

OpenStudy (anonymous):

ohhh i see

OpenStudy (anonymous):

so lemme work this all out

OpenStudy (zale101):

Okay, thanks for participating :)

OpenStudy (jhannybean):

just use the complete the square method :P

OpenStudy (anonymous):

wait dont confuse me now XD

OpenStudy (jhannybean):

Oh no, check this later after you have solved it on your own. Okay?

OpenStudy (anonymous):

zale when i take the half and square it where does that go? im still new to all this XD

OpenStudy (zale101):

when you complete the square, you'd get 4 4/2=2 2^2=4 add and subtract 4 into the parenthesis and factor the (x^2+4x+4) \(\large (x^2+4x+4)-4=y-5\)

OpenStudy (zale101):

\((x^2+4x+4)\) is in perfect trinomial, you know the factor for it?

OpenStudy (anonymous):

(x+4)(x+1) i think

OpenStudy (zale101):

Perfect trinomial, meaning that (x+2)(x+2) is the factor, \((x+2)^2\). \(\Large(x+2)^2-4=y-5\)

OpenStudy (anonymous):

oh ok i see

OpenStudy (zale101):

What is the next step? We already have h \(\Large (x-\color{red} h)^2+\color{blue}k=y\) where h is -2 \(\Large(x-(\color{red}{-2}))^2-4=y-5\) how do you get k?

OpenStudy (anonymous):

add 5 to both sides?

OpenStudy (zale101):

yes

OpenStudy (anonymous):

got it thank you so much!

OpenStudy (mchilds15):

Lol Zale is such a smarty pants. Good job :)

OpenStudy (zale101):

No problem, but you know why i got -2? because we had (x+2)^2 and h is written in the vertex form as (x-h)^2 that means that 2 is negative \((x-(-2))^2\)

OpenStudy (anonymous):

got it :) thanks zale

OpenStudy (zale101):

you're welcome~